# Climb: Class-imbalanced Learning Benchmark on Tabular Data

### v1

@@ A.3 Additional Results
Detailed main results on each dataset.
@@
Due to space constraints, in the main results (Table 2), we reported the average scores and rankings for each metric by grouping the 73 datasets into four categories based on their imbalance levels. Here, we provide the complete results for each method on each individual dataset. Specifically, AUPRC, F1-score, and Balanced Accuracy results are reported in Tables 5, 6, and 7, respectively. Additionally, Table 8 presents the runtime of each method across different datasets. The dataset ordering in these tables follows the order defined in Table 3. We used color coding similar to Table 2 (i.e., blue/red represents better/worse than no-balancing, deeper color means larger difference) for better visual clarity.
@@
Table 5: Detailed full results on each dataset on AUPRC (×$\times 10^{-2}$).
@@
Dataset Base Undersample Cleaning Oversample Undersample Ensemble Oversample Ensemble Cost-senstive

### v2

@@ C.1 Results with Additional CIL methods
C.3 Discussion on the Self-paced Ensemble
Among all the evaluated CIL methods, Self-paced Ensemble (SPE) stands out as the most consistent top performer. Its advantage can be attributed to a few complementary factors.
  • Hard example mining: SPE retains difficult-to-classify samples during undersampling, which improves decision boundaries.
  • Noise robustness: The hardness harmonization mechanism balances informativeness and noise, avoiding the inclusion of overly noisy samples.
  • Self-paced learning: Inspired by curriculum learning, SPE introduces samples progressively from easy to hard, which stabilizes training.
  • Efficiency: As an undersampling-based ensemble, SPE trains on fewer samples per model, making it more efficient than oversampling or boosting strategies.
C.4 Comparison with BAF Benchmark
We also compare CLIMB with the BAF benchmark (Jesus et al., 2022). Both address imbalance, but with different goals and setups. CLIMB evaluates CIL methods on 73 real-world datasets with natural imbalance, using repeated cross-validation and standard metrics such as AUPRC, Macro-F1, and Balanced Accuracy. BAF instead focuses on fairness under distributional bias and temporal shift in a single fraud detection task, using CTGAN-generated synthetic data, temporal splits, and fairness metrics like TPR@5%FPR and FPR ratio. Thus, CLIMB offers a broad benchmark for CIL effectiveness, while BAF targets fairness in a specific application.
C.5 Detailed main results on each dataset.
@@ C.4 Comparison with BAF Benchmark
Due to space constraints, in the main results (Table 2), we reported the average scores and rankings for each metric by grouping the 73 datasets into four categories based on their imbalance levels. Here, we provide the complete results for each method on each individual dataset. Specifically, AUPRC, F1-score, and Balanced Accuracy results are reported in Tables 6, 7, and 8, respectively. Additionally, Table 9 presents the runtime of each method across different datasets. The dataset ordering in these tables follows the order defined in Table 3. We used color coding similar to Table 2 (i.e., blue represents better than no balancing, and red represents worse than no balancing, with deeper colors indicating larger differences) for improved clarity.
Dataset-level Analysis.
Although the overall conclusions of CLIMB are robust across datasets, a few cases deviate from the general trends. We intentionally phrased our main takeaways to avoid overgeneralization, and here we highlight notable examples to provide additional context:
  • Undersampling ensembles on extremely imbalanced datasets (e.g., dis, satellite): Random undersampling based ensembles such as Balanced Random Forest (BRF), EasyEnsemble, and UBS can fail when the imbalance ratio is very severe. These methods discard most majority class samples, which results in insufficient training information and weak generalization. In contrast, approaches like Self-paced Ensemble (SPE) and BalanceCascade (BC) are more robust because they explicitly retain informative samples through hard example mining.
  • Cleaning based methods on long-tailed multiclass datasets (e.g., user-knowledge, allbp): Cleaning based methods such as Tomek Links, ENN, and RENN often underperform in long-tailed multiclass scenarios. Since multiple minority classes can be close to majority classes in feature space, these cleaning procedures tend to over remove minority samples. This reduces the model’s ability to learn rare class patterns and leads to degraded performance.
These exceptions are limited in scope and do not alter the overall conclusions of our study. Instead, they illustrate the importance of understanding dataset specific characteristics when selecting and applying CIL methods in practice.
@@
Table 6: Detailed full results on each dataset on AUPRC ($\times 10^{-2}$).
@@
Dataset Base Undersample Cleaning Oversample Hybrid Undersample Ensemble Oversample Ensemble Cost-Sensitive GBDTs